OpenStudy (anonymous):

The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line?

OpenStudy (anonymous):

y =( –(5/2)x) – 5 y = (2/5)x) + 2 y =(5/2)x)– 2 y = –(2/5)x) – 2

OpenStudy (anonymous):



OpenStudy (anonymous):

can you find the slope of this line?

OpenStudy (anonymous):

-2/5

OpenStudy (anonymous):

yep ^_^ sooo, we know m=-2/5 and we know the 'slope intercept' form of a line is y = mx+b y = (-2/5)x +b so all we need now is 'b'

OpenStudy (anonymous):

'b' is the number where the line intersects the y-axis, do you know what that number is?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

yep ^_^ so we had y = (-2/5)x +b and now know b = -2 so replace b with -2 in our equation and you got the answer ^_^